OpenStudy (anonymous):

Write the recursive formula for the geometric sequence.a^1 = -2 a^2 = 8 a^3 = -32

OpenStudy (helder_edwin):

\[ \large a_1=-2=-2\cdot(-4)^0 \] \[ \large a_2=8=(-2)(-4)=a_1\cdot(-4) \] \[ \large a_3=-32=8(-4)=a_2\cdot(-4)=a_1\cdot(-4)(-4)=a_1\cdot(-4)^2 \]